### v2

Abstract
Melanoma is a malignant tumor originating from skin cell lesions. Accurate and efficient segmentation of skin lesions is essential for quantitative medical analysis but remains a challenge. To address this, we propose ScaleFusionNet, a segmentation model that integrates the Cross-Attention Transformer Module (CATM) and AdaptiveFusionBlock to enhance feature extraction and fusion. The model employs a hybrid architecture encoder that effectively captures both local and global features. We introduce CATM, which utilizes Swin Transformer Blocks and Cross Attention Fusion (CAF) to adaptively refine encoder-decoder feature fusion, reducing semantic gaps and improving segmentation accuracy. Additionally, the AdaptiveFusionBlock is improved by integrating adaptive multi-scale fusion, where Swin Transformer-based attention complements deformable convolution-based multi-scale feature extraction. This enhancement refines the boundaries of the lesion and preserves fine-grained details. ScaleFusionNet achieves Dice scores of 92. 94% and 91. 65% in the ISIC-2016 and ISIC-2018 datasets, respectively, demonstrating its effectiveness in skin lesion analysis. Our code implementation is publicly available at GitHub.
keywords:
Transformer, Skin Lesion, Image Segmentation, Information fusion, feature enhancement
1 Introduction
The incidence of melanoma has risen significantly in recent decades, driven by increasing environmental pollution and ultraviolet radiation [1]. This trend has drawn significant attention from the global medical community. Early detection is essential for improving treatment outcomes and patient survival. Traditional diagnostic methods, such as clinical observation and tissue biopsy, are limited by subjectivity and invasiveness, which make them unsuitable for large-scale screening. Medical image segmentation offers a noninvasive and high-precision alternative that gives clinicians with more detailed and accurate information. This technology has the potential to greatly enhance early detection and treatment of skin cancer.
With the rapid advancement of deep learning, convolutional neural network (CNN)-based medical image segmentation models have gained prominence, achieving notable success in tasks such as skin lesion segmentation. Among these, the U-Net model, introduced by Ronneberger et al. [2], stands out as a pioneering framework. U-Net has demonstrated remarkable performance in medical image segmentation and has established a distinctive U-shaped architectural paradigm. However, due to the inherent limitations of convolutional operations, CNN-based methods often struggle to model global contextual information effectively. This limitation becomes particularly evident in medical image segmentation tasks with significant inter-sample variability, such as skin lesion segmentation. To address this challenge, researchers have explored various strategies, including the use of large kernel convolutions, dilated convolutions, and other techniques aimed at expanding receptive fields [3, 4, 5, 6, 7]. For example, Hu et al. [8] employed self-attention to enhance readvantageelds, while Tang et al. [9] proposed a model that took adavantage of large convolutional kernels and skip fusion to achieve promising results in tasks such as breast nodule ultrasound image segmentation. Inspired by the ConvNeXt model [10], Han et al. [11] developed a medical image segmentation method based on large kernel convolutions, successfully applying it to tasks such as retinal vessel segmentation. Despite these advancements, simply increasing the size of convolutional kernels may not fully resolve the challenge of modeling global features, as the fundamental constraints of the receptive field remain.
Recently, Transformer [12], which serves as a feature extraction technique with powerful global modeling capabilities, has achieved notable success in both the natural language processing and computer vision domains. Cai et al. [13] unveiled the BiADATU‐Net that combined Transformer and feature adaptation modules, which resulted in promising outcomes in several publicly accessible skin lesion segmentation datasets. Zhang et al. [14] developed DAE-Former, a pure Transformer U-shaped medical image segmentation model, harnessing efficient Transformers steered by dual attention, similar to Swin-Unet [15], and exhibited commendable performance in diverse image segmentation datasets, including ISIC-2018 [16]. However, because Vision Transformers can only output single-scale feature representations, they lack the ability to capture multi-scale information in two-dimensional images [17, 18]. As a result, Transformer-based medical image segmentation models may struggle to seamlessly integrate multi-scale information, leading to insufficient attention to lesion regions and incomplete decoding of feature details.
Additionally, there exists a significant issue in medical image segmentation models based on the U-Net design architecture. Although the skip connections in U-Net can transmit multi-scale information between different stages to the decoder, a semantic gap issue may arise when there is a considerable semantic difference between the encoder and decoder features. To address this, some research endeavors have attempted to mitigate this issue by improving skip connections. For instance, UNet++ [19] and MISSFormer [20] aim to achieve the fusion of multi-scale information between different stages through dense skip connections and contextual bridges. Nevertheless, this paper argues that the different-sized feature maps transmitted through skip connections represent macroscopic multi-scale information that is easily observable to the naked eye, and such methods have limited effectiveness in enhancing the model’s ability to integrate multi-scale information. Particularly in the skin lesion segmentation task, the lesion edges are often irregular, with colors gradually fading from the center, and the progressive compression of feature maps leads to the loss of fine details, retaining only macro-level focus. This can, to some extent, affect the model’s performance.
To address the challenges in medical image segmentation, this paper proposes ScaleFusionNet, a model that integrates an AdaptiveFusionBlock and a Cross-Attention Transformer Module (CATM) for enhanced feature extraction and fusion. ScaleFusionNet employs a hierarchical Swin Transformer-based encoder, where patch embedding and Swin Transformer blocks [21] extract multi-scale features. The decoder utilizes AdaptiveFusionBlocks, which combine deformable convolutions and attention mechanisms to refine feature integration and improve lesion boundary preservation. To bridge the semantic gap between encoder and decoder features, CATM leverages cross-attention, allowing high-level decoder features to guide low-level skip connections. Experimental results demonstrate that ScaleFusionNet achieves highly competitive performance in skin lesion segmentation. The key contributions of this work are as follows:
  • Proposed an advanced medical image segmentation network named ScaleFusionNet, based on a hybrid architecture combining CNNs and Transformers.
  • Introduced AdaptiveFusionBlock that integrates Swin Transformer-based feature extraction and adaptive multi-scale fusion.
  • Developed the Cross-Attention Transformer Module (CATM) to effectively reduce the semantic gap and enhance encoder-decoder feature interaction.
  • Achieved highly competitive results on ISIC-2016 and ISIC-2018, demonstrating superior performance in skin lesion segmentation.
2 Related Work
2.1 CNNs for Medical Image Segmentation
In recent years, Convolutional Neural Networks (CNNs) have achieved notable success in various domains due to their powerful feature extraction capabilities. This success is particularly evident in the field of medical image segmentation. In 2015, Ronneberger et al. introduced U-Net, a CNN-based model designed specifically for medical image segmentation, which has become foundational in this field. To address the semantic gap between the encoder and decoder, Zhou et al. proposed UNet++, which introduces nested dense skip connections. Li et al. [22] developed an H-DenseUNet, a U-shaped model that enhances intra-slice and inter-slice representations through hybrid dense connections, demonstrating effective performance in liver tumor segmentation tasks. Saqib et al. [23] proposed multi-scaled architecture using separable convolution for brain tumor segmentation. Oktay et al. [24] proposed Attention U-Net, a model that learns to suppress irrelevant regions using attention gates, effectively mitigating the semantic gap between encoder and decoder. Furthermore, UNet3+ [25] advanced the skip connection design by incorporating full-scale skip connections and deep supervision, achieving improvements in segmentation tasks. Despite the success of U-Net and its variants, CNN-based models are limited by their inability to capture long-range dependencies. This limitation arises due to the inherent nature of convolution operations, which struggle to model global contextual information. Moreover, dense skip connections based on simple summation offer limited solutions for addressing the semantic gap, especially in tasks where fine-grained detail and global context are crucial, such as skin lesion segmentation.
2.2 Transformers for Medical Image Segmentation
Transformers, which are adept at capturing long-range dependencies, offer an effective alternative to CNNs. The Vision Transformer [26], the first application of Transformer to computer vision, partitions input images into a sequence of patches for embedding and encoding using Transformer blocks. This innovative approach inspired Transformer-based U-shaped models for medical image segmentation tasks. TransUnet [27] integrated Transformer blocks with U-Net, leveraging their global feature modeling capability. In parallel, Swin-Unet [15] drew on the Swin Transformer to propose a fully Transformer-based method that applies Swin Transformer blocks to both the encoder and decoder. For 3D medical image segmentation, nnFormer [28] combines local and global attention for multi-organ segmentation, demonstrating impressive performance. However, Transformer models, such as TransUnet, face challenges due to the increased number of parameters and computational complexity, and the Swin Transformer’s shifting window operation can introduce jagged boundaries in certain cases. Furthermore, Transformers inherently operate on single-scale outputs, limiting their ability to fully utilize multi-scale information, which hampers performance, especially in segmentation tasks requiring precise boundary delineation.
2.3 Deformable Convolution Network
The Deformable Convolutional Network (DCN) [29], proposed by Dai et al. in 2017, is an extension of traditional convolutional networks designed to improve the adaptability of convolutional kernels to object deformations. DCNs introduce learned offsets to control the sampling positions of kernels on input feature maps. This dynamic adjustment allows the model to better capture the features of targets with varying shapes and positions. Compared to traditional convolution, deformable convolutions offer more flexibility in capturing features of objects with diverse shapes and scales, which is crucial in medical image segmentation. DCNs have proven effective in addressing the complexities of various segmentation tasks, where object deformations are a significant concern [30].
2.4 Multi-Scale and Hybrid Architectures
In the pursuit of more accurate medical image segmentation, particularly for skin lesion segmentation, ScaleFusionNet integrates multi-scale feature extraction with a hybrid Transformer-CNN architecture to enhance segmentation performance. Traditional CNN-based models like U-Net struggle with fine-grained details and global context, prompting the need for improved architectures. ScaleFusionNet addresses these limitations by incorporating AdaptiveFusionBlocks, which leverage deformable convolutions [29] for multi-scale feature extraction, refining lesion boundaries while preserving fine details. Additionally, the Cross-Attention Transformer Module enhances encoder-decoder feature fusion, reducing semantic gaps through guided attention mechanisms. By combining Swin Transformer Blocks [21] for global context and adaptive multi-scale fusion for local detail refinement, ScaleFusionNet achieves superior segmentation accuracy, demonstrating strong generalization in challenging skin lesion segmentation tasks.
3 Methods
The architecture of ScaleFusionNet, as illustrated in Figure 1, follows the U-Net design and consists of three primary components: the encoder, CATM, and the AdaptiveFusionBlock. The encoder employs a hybrid approach that integrates convolutional layers and Swin Transformer Blocks to capture both local and global features effectively. By combining convolutional locality with self-attention mechanisms, the encoder enhances feature representation, ensuring robust extraction of hierarchical information. The CATM is introduced at skip connections to refine encoder-decoder feature fusion. It utilizes Swin Transformer Blocks and Cross Attention Fusion (CAF) to mitigate the semantic gap and align hierarchical features dynamically. By leveraging cross-attention mechanisms, CATM enhances the integration of skip connection features with decoder information, ensuring improved contextual understanding. The AdaptiveFusionBlock further enhances multi-scale feature extraction and fusion by integrating deformable convolutions with Swin Transformer-based attention mechanisms. This fusion process refines lesion boundaries and preserves fine-grained details, which is essential for accurate segmentation. Given an input image $I\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ denote the height, width, and number of channels, respectively, the encoder progressively extracts multi-scale features. The CATM is positioned at skip connection points, receiving features $X_{\text{Skip}}$ from the encoder and $X_{\text{Decoder}}$ from the lower-level decoder. The extracted features are dynamically refined using cross-attention mechanisms, allowing precise alignment and feature enhancement. The AdaptiveFusionBlock processes these refined features by fusing multi-scale information through deformable convolutions and Swin Transformer-based attention. The final segmentation result is obtained after feature fusion and upsampling operations in the decoder, ensuring high-resolution and well-defined segmentation masks.
Figure 1: Architecture of ScaleFusionNet.
3.1 CATM
In U-Net, skip connections serve as a means of information supplementation. During the encoding process, continuous compression of feature maps leads to a significant loss of spatial details. Using low-level semantic features from the encoder to supplement the decoder’s feature restoration is an effective strategy. However, a fundamental issue remains: the semantic gap between encoder and decoder features. Directly concatenating features at different semantic levels can result in performance degradation due to this misalignment.
To address this, we introduce CATM to refine encoder-decoder feature fusion. Unlike conventional skip connections, CATM employs Swin Transformer Blocks and Cross Attention Fusion to adaptively align hierarchical features. By leveraging self-attention and cross-attention mechanisms, CATM ensures the effective transfer of relevant spatial and contextual information across the network.
As shown in Figure 2, given the encoder features $X_{\text{Skip}}$ and the decoder features $X_{\text{Decoder}}$, CATM dynamically refines the feature alignment using a learnable attention mechanism. This process is formulated as follows:
$Q,K,V=\text{SwinTransformerBlock}(X_{\text{Decoder}})$
$X_{\text{CATM}}=\text{CAF}(X_{\text{Skip}},Q,K,V)$
Here, the Swin Transformer Block extracts query, key, and value representations from the decoder features, while Cross Attention Fusion (CAF) integrates these with encoder features $X_{\text{Skip}}$, reducing the semantic gap. The refined features $X_{\text{CATM}}$ preserve both fine-grained details and high-level contextual information.
Figure 2: Schematic diagram of CATM structure.
Algorithm 1 Cross-Attention Transformer Module (CATM) for ScaleFusionNet
1:$X_{\text{Decoder}},X_{\text{Skip}}\in\mathbb{R}^{H\times W\times C}$
2:$X_{\text{CATM}}\in\mathbb{R}^{H\times W\times C}$
3:$Q,K,V=\text{SwinTransformerBlock}(X_{\text{Decoder}})$triangleright Obtain Query, Key, and Value
4:$X^{\prime}_{\text{Skip}}=\text{CAF}(X_{\text{Skip}},Q,K,V)$triangleright Feature Alignment
5:$X_{\text{CATM}}=\text{SharedSA}(X^{\prime}_{\text{Skip}})$triangleright Refined Feature Fusion
6:return $X_{\text{CATM}}$triangleright Return
3.2 Adaptive Fusion Block
The decoder plays a critical role in feature decompression and mask generation for medical image segmentation. A key challenge is accurately restoring boundary details and enhancing attention toward target regions. Conventional decoder designs, whether convolution-based or Transformer-based, often struggle to effectively capture fine-scale information, leading to imprecise lesion localization and segmentation.
To address these limitations, we introduce the AdaptiveFusionBlock, which integrates adaptive multi-scale feature fusion to refine segmentation. This block leverages Swin Transformer-based attention and deformable convolution-based multi-scale feature extraction, allowing the model to adaptively capture both local and global contextual information.
As shown in Figure 2, given the features from the encoder $X_{\text{Skip}}$ and decoder $X_{\text{Decoder}}$, the AdaptiveFusionBlock first concatenates these features along the channel dimension:
$X=\text{Concat}(X_{\text{Decoder}},X_{\text{Skip}})$
The input feature $X$ is then processed through a hybrid multi-scale extractor, which consists of Swin Transformer Blocks for capturing long-range dependencies and deformable convolutions for learning spatially adaptive features. The multi-scale feature extraction process is formulated as:
$X^{\prime}_{\text{Swin}}=\text{SwinTransformer}(X)$
$X^{\prime}_{\text{Deform}}=\text{DeformConv}(X)$
The outputs from both branches are fused to generate the refined feature representation:
$X_{\text{AFB}}=X+\text{Concat}(X^{\prime}_{\text{Swin}},X^{\prime}_{\text{Deform}})$
This fusion mechanism enhances lesion boundary refinement while preserving fine-grained details. Finally, the fused features are passed through an upsampling module to ensure effective decoding.
By integrating the AdaptiveFusionBlock into ScaleFusionNet, we achieve improved segmentation performance by dynamically adapting feature extraction and fusion strategies, leading to better lesion delineation and generalization across datasets.
Figure 3: Schematic diagram of AdaptiveFusionBlock structure.
Algorithm 2 AdaptiveFusionBlock for ScaleFusionNet
1:$X_{\text{Decoder}},X_{\text{Skip}}\in\mathbb{R}^{H\times W\times C}$
2:$X_{\text{Out}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C^{\prime}}$
3:$X=\text{Concat}(X_{\text{Decoder}},X_{\text{Skip}})$triangleright Concatenate features
4:$X^{\prime}=\text{Conv2D}_{1\times 1}(X)$triangleright Apply 1x1 convolution
5:$X_{\text{Swin}}^{\prime}=\text{SwinTransformer}(X^{\prime})$triangleright Extract features using Swin Transformer
6:$X_{\text{Deform}}^{\prime}=\text{DeformConv}(X^{\prime})$triangleright Extract features using deformable convolution
7:$X^{\prime\prime}=X^{\prime}+\text{Concat}(X_{\text{Swin}}^{\prime},X_{\text{Deform}}^{\prime})$triangleright Fuse features
8:$X^{\prime\prime\prime}=\text{UpConv}(X^{\prime\prime})$triangleright Upsample features
9:$X_{\text{Out}}=X^{\prime\prime\prime}+\text{LayerScale}(X^{\prime\prime\prime})$triangleright Apply layer scaling
10:return $X_{\text{Out}}$triangleright Return the output features
4 Experiments
4.1 Datasets
1) ISIC-2016
This dataset is derived from the Skin Lesion Analysis Towards Melanoma Detection challenge in 2016, comprising 1250 images meticulously annotated by professional experts with gold standard labels [31]. Among these, 900 images are designated as training data, while 350 images are allocated for validation.
2) ISIC-2018
The ISIC-2018 dataset, also collected by ISIC in 2018, consists of 2594 images and corresponding labels. The resolutions of the images range from $720\times 540$ to $6708\times 4439$ pixels [16]. Among these, 2594 images are randomly divided into training, validation, and test sets at a ratio of 8:1:1.
4.2 Evaluation metrics
In this paper, the main evaluation metrics employed include the Dice coefficient (also known as Dice similarity score, DSC), Intersection over Union (IoU).
4.2.1 Dice Coefficient (DSC)
The Dice coefficient measures the overlap between the predicted segmentation mask and the ground truth mask. It is defined as:
$DSC=\frac{2|A\cap B|}{|A|+|B|}$
where $A$ is the predicted mask and $B$ is the ground truth mask. A higher Dice score indicates better segmentation performance.
4.2.2 Intersection over Union (IoU)
Intersection over Union evaluates segmentation accuracy by computing the ratio of the intersection to the union between the predicted and actual masks:
$IoU=\frac{|A\cap B|}{|A\cup B|}$
4.3 Implement details
All experiments in this paper are conducted using the PyTorch 1.12.0 framework. The experiments are carried out on a computer equipped with the Ubuntu 18.04 operating system, an Intel Core i9-13900K CPU, an Nvidia RTX 4060 GPU, and a 1TB solid-state drive. For all experiments involving ScaleFusionNet, the AdamW optimizer is utilized with a learning rate and weight decay set to 1e-4. We used a combination of BCE loss and IOU loss to form our loss function, along with random rotation and random flipping for data augmentation. For comparison purposes, we referenced the experimental results disclosed in relevant papers for similar methods. For outstanding models that did not perform skin lesion segmentation tasks, we retrained them using their publicly available official execution code. To ensure fairness, we kept parameters that do not affect model learning capacity, such as epochs and batch size, consistent with ScaleFusionNet, setting epochs to 200 and batch size to 8. The experiments focused on the ISIC-2016 and ISIC-2018 datasets.
This experimental setup ensures a robust and fair evaluation of ScaleFusionNet’s performance, leveraging state-of-the-art hardware and software configurations to achieve accurate and reproducible results. The use of a combined loss function and data augmentation techniques further enhances the model’s ability to generalize and perform well on diverse skin lesion segmentation tasks.
5 Experimental Results
5.1 Results on ISIC-2016
To better compare with current mainstream models, 11 prominent models were selected for comparison with the proposed ScaleFusionNet. Table 1 shows that ScaleFusionNet achieved a DSC score of 92.94% and an IOU score of 87.35% on the ISIC-2016 dataset, which are the average results of five repeated experiments, demonstrating outstanding performance. Compared to the Swin-Unet model, ScaleFusionNet improved by 2.82% in the DSC metric and 4.14% in the IOU metric. When compared to MISSFormer, ScaleFusionNet exhibited enhancements of 2.48% and 3.43% in the DSC and IOU metrics, respectively. Against the D-LKA model, ScaleFusionNet still showed improvements of 0.11% and 0.20% in the DSC and IOU metrics, respectively. This indicates that ScaleFusionNet more accurate than D-LKA to detect the refined boundary of skin lesions. Compared to methods like U-Net, although ScaleFusionNet employs a more complex architecture to address potential issues in U-shaped medical image segmentation models, we find this approach justified given the 5.13% performance improvement and the reduced computational load. In clinical applications, a faster and lighter model can support a broader range of compatible use cases, which is crucial for hospitals or organizations with limited computational resources. Concurrently, to enhance the assessment of model performance, we select 7 high-performing models for qualitative scrutiny of experimental outcomes, elucidating the differences among them. The red areas in the illustrations represent ground-truth labels meticulously annotated by experts, reflecting the diagnostic preferences of clinical doctors in real scenarios. A larger red area indicates lower model accuracy and poorer discrimination of affected regions. In contrast, the green areas represent the predicted labels obtained during the model testing phase. A larger green area suggests that the model has mistakenly segmented healthy skin, which could mislead doctors into treating non-affected areas, especially with destructive procedures like lasers or cryotherapy. The yellow areas indicate the overlap between the predicted and ground truth labels; a larger yellow area signifies a more accurate identification of the region of the lesion by the model. In summary, from the perspective of clinical diagnosis and treatment, smaller green and red areas and a larger yellow area indicate better model performance, enabling more effective segmentation of skin lesions to help in diagnosis and treatment decisions.
The results of the qualitative analysis of the 7 models in the ISIC-2016 dataset are shown in Figure 5(a). From the first two rows, it is evident that ScaleFusionNet exhibits a broader yellow region compared to the D-LKA model, indicating that ScaleFusionNet is better at identifying affected areas. Furthermore, ScaleFusionNet’s predictions show fewer green and red regions, which reduces the likelihood of misdiagnosis and missed diagnoses. This difference is even more pronounced in the latter two rows. Although TransFuse has the largest yellow region, it also shows the largest green area, indicating a misdiagnosis of healthy regions. Although it covers areas of injury, this misdiagnosis can have serious implications in the clinical diagnosis. In comparison to D-LKA, ScaleFusionNet has a similarly sized yellow region but with a smaller misdiagnosis area, aligning better with clinical diagnostic needs.
These results highlight ScaleFusionNet’s superior ability to accurately segment skin lesions while minimizing errors, making it a highly effective tool for clinical applications. Its combination of high accuracy, and strong generalization ability positions it as a leading solution for skin lesion segmentation tasks. The model’s ability to preserve fine-grained details and refine the limits of the injury further underscores its potential to improve melanoma diagnosis and treatment in real-world clinical settings.
Table 1: Performance Comparison on ISIC-2016 Dataset
Methods DSC IOU
U-Net [2] 87.81 80.25
Att-Unet [24] 87.43 79.70
nnU-Net [32] 90.45 84.52
SwinUNet [15] 90.12 83.21
MISSFormer [20] 90.46 83.92
DAEFormer [14] 91.19 85.40
HiFormer [33] 91.48 85.15
TransFuse [34] 92.03 86.19
D-LKA [35] 92.83 87.15
SU-Net [36] 92.33 86.58
ScaleFusionNet (Ours) 92.94 87.35
5.2 Results on ISIC-2018
In the experiments conducted on the ISIC-2018 dataset, 11 mainstream medical image segmentation models were selected for comparison with ScaleFusionNet, and more relevant parameters were disclosed here. The reason for the incomplete consistency in model selection lies in the absence of relevant results and code publication in the papers of certain models, or the inability to reproduce the code provided by them. The quantitative analysis results of the ISIC-2018 comparison experiments are presented in Table 2.
From Table 2, it can be observed that ScaleFusionNet continues to exhibit competitive results compared to the other 20 medical image segmentation methods, achieving the best results in terms of the DSC metric on ISIC-2018, with most other metrics ranking in the top two. Compared to TransFuse, ScaleFusionNet showed a 0.57% improvement in the DSC metric while maintaining a consistent level of IOU as 0.92%. Compared to D-LKA, ScaleFusionNet exhibited a 0.01% improvement in the DSC metric, although it slightly lagged behind D-LKA in the IOU metric. However, ScaleFusionNet has competitive results with other methods which shows strong performance in skin lesion segmentation.
Table 2: Performance Comparison on ISIC-2018 Dataset
Methods DSC IOU
U-Net [2] 85.45 77.33
Att-Unet [24] 85.66 77.64
nnU-Net [32] 89.03 82.02
Swin-Unet [15] 89.31 82.14
MISSFormer [20] 89.44 82.41
DAEFormer [14] 89.89 83.21
HiFormer [33] 90.55 83.81
TransFuse [34] 91.08 84.65
D-LKA [35] 91.64 85.64
SU-Net [36] 90.90 84.49
ScaleFusionNet (Ours) 91.65 85.57
Figure  4 visually demonstrates the results of various segmentation models in the ISIC 2016 and 2018 datasets, focusing on both DSC and IOU metrics. As observed, ScaleFusionNet consistently outperforms other models, particularly in the ISIC 2016 dataset, where it achieves the highest DSC score. The blue line for ISIC 2016 DSC shows a steady increase, with ScaleFusionNet clearly marked as the top performer. Similarly, for ISIC 2018 DSC, ScaleFusionNet, indicated by the green line, again achieves a leading position, closely followed by models such as D-LKA and TransFuse. The dashed red and yellow lines represent the IOU scores, where ScaleFusionNet excels in the 2016 dataset, further demonstrating its dominance in multiple evaluation metrics. Overall, the graph illustrates the effectiveness of ScaleFusionNet in both DSC and IOU metrics for both datasets, demonstrating it to be a robust model for skin lesion segmentation tasks.
Figure 4: Performance comparison of segmentation models on ISIC-2016 and ISIC-2018 datasets. The figure shows DSC (circles) and IOU (squares) scores for various models, with ScaleFusionNet outperforming other models in both metrics across both datasets.
The qualitative analysis results of ISIC-2018 are illustrated in Figure 5(b). Visual inspection of the results in the first two rows reveals that, compared to D-LKA, ScaleFusionNet has a larger yellow region and a smaller red region, indicating higher prediction accuracy, even though it performs slightly worse on the IOU metric. In comparison to PVT-GCASCADE, while the yellow regions are nearly identical in size, ScaleFusionNet exhibits a smaller range of green areas, demonstrating its superior ability to identify skin lesion regions and reduce the likelihood of misdiagnosis. Similarly, the results in the last two rows further highlight the overall superior performance of ScaleFusionNet.
Based on the experiments using the ISIC-2018 dataset, the D-LKA model and ScaleFusionNet achieved the first and second best performances in terms of the DSC and IOU metrics, respectively. In terms of overall performance, ScaleFusionNet demonstrates better accuracy and boundary fitting. The qualitative analysis of the results, as illustrated in Figure 5(b), shows that ScaleFusionNet has a smaller green region compared to D-LKA, indicating a reduced likelihood of misdiagnosis. This further underscores ScaleFusionNet’s ability to accurately segment skin lesions while minimizing errors. These results highlight ScaleFusionNet’s superior performance on the ISIC-2018 dataset, achieving high accuracy in lesion segmentation while maintaining efficiency in terms of parameter count and computational complexity. Its ability to reduce misdiagnosis and improve lesion boundary delineation makes it a highly effective tool for skin lesion analysis, particularly in clinical applications where precision and efficiency are critical.
Figure 5: Qualitative analysis of 2016 and 2018 ISIC datasets.
5.3 Ablation Study
To validate the effectiveness of the proposed ScaleFusionNet, structural ablation study is conducted on the ISIC-2016 dataset, with the DSC used as the primary evaluation metric. The data from the structural ablation studies are presented in Table 3. Method 0 represents the segmentation results obtained using only the hybrid architecture, Method 1 represents the experimental results obtained using both the hybrid architecture and the Cross-Attention Transformer Module (CATM), and Method 2 represents the experimental results obtained using only the hybrid architecture and the AdaptiveFusionBlock. From Table 3, it can be observed that using only the hybrid architecture results in a DSC of 91.76% on ISIC-2016, which is still superior to most models compared in Table 1. Method 1, which incorporates the CATM to enhance the skip connection features on top of the hybrid architecture, achieves a DSC of 92.54% on ISIC-2016. This performance surpasses that of the SU-Net method at 92.33%, albeit slightly weaker than the D-LKA method. Method 2, which introduces only the AdaptiveFusionBlock on top of the hybrid architecture, achieves a DSC performance of 92.62%. Finally, ScaleFusionNet, which combines the hybrid architecture, CATM, and AdaptiveFusionBlock, achieves a DSC of 92.94% on ISIC-2016. The excellent performance of ScaleFusionNet is propelled by the integration of these three proposed improvement methods. Furthermore, to intuitively observe the role of each structure, the features of Stage 0 and the corresponding modules within the same layer are visualized. As indicated in Figure 6, it is visually apparent that the attention intensity of the target area is significantly enhanced after passing through the CATM following the output of the hybrid architecture. Additionally, from the feature maps of the four multi-scale branches, it can be observed that each feature map highlights different attention areas. This underscores the emphasis on micro-scale multi-resolution in this study. The micro-scale multi-resolution enables the features passed through the AdaptiveFusionBlock to focus highly on the target area and exhibit good fitting to the target boundaries.
Table 3: Structural ablation study of ScaleFusionNet on ISIC-2016.
[…]
0 91.76
1 92.54
2 92.62
ScaleFusionNet 92.94
These ablation experiments demonstrate the critical contributions of the CATM and AdaptiveFusionBlock to the performance of ScaleFusionNet, as well as the importance of optimizing the encoder’s architecture. The results highlight the model’s ability to achieve high accuracy in skin lesion segmentation while maintaining an efficient and balanced architecture.
Figure 6: Features visualization of different layers in ScaleFusionNet.
6 CONCLUSION
This study presents a medical image segmentation model called ScaleFusionNet, which incorporates the Cross-Attention Transformer Module (CATM) and AdaptiveFusionBlock. The model utilizes deep learning algorithms to learn and extract complex features from medical images. Specifically, the proposed AdaptiveFusionBlock and CATM significantly enhance performance in skin cancer segmentation tasks. Experiments conducted on publicly available datasets demonstrate that this model achieves competitive results, strongly supporting advances in skin cancer medical image segmentation. These innovative approaches positively impact diagnostic accuracy, guide treatment decisions, and promote further research in the field. However, there are still some issues within this study. While ScaleFusionNet maintains a higher accuracy, its computational complexity is higher compared to some methods. This is primarily due to the integration of multi-scale feature extraction and cross-attention mechanisms, where contributions from different receptive fields may vary. A potential solution could involve optimizing feature allocation in a manner similar to self-selective routing. Additionally, the application of deformable convolutions may exacerbate this issue, making it worthwhile to explore simplifications in deformable operations in future research.
From a clinical application perspective, a reliable medical image segmentation method must not only provide high-quality segmentation results but also deliver corresponding uncertainty metrics. ScaleFusionNet represents a significant step forward in skin lesion segmentation, leveraging adaptive multi-scale fusion and cross-attention mechanisms to achieve high accuracy and robustness. However, future work should address computational efficiency and incorporate uncertainty quantification to further enhance its clinical applicability and reliability. By aligning more closely with clinical needs, such as accounting for the severity and variability of lesions, ScaleFusionNet and its successors can better support dermatologists in diagnosing and treating skin conditions effectively.
[…]
Acknowledgements
This work was supported in part by the UK EPSRC under Grant EP/P009727/2, and the Leverhulme Trust under Grant RF-2019-492. The authors also express their appreciation to Taif University, Saudi Arabia, for supporting this work through project number (TU-DSPP-2024-229).
Declarations
Availability of Data and Materials: The datasets analyzed in the current study are available in the 2016 and 2018 repositories at the following link: https://challenge.isic-archive.com/data/.
Conflict of interest: The authors declare no competing interests.